# HateDebias: On the Diversity and Variability of Hate Speech Debiasing

### v1

@@ 4 Conclusion
In addition to conducting experiments on the BERT model, we further perform our method with other pre-trained language models Deberta-v3-base [58] and RoBERTa-base [59]. The experimental results are shown in the Table A3. Experimental results verify the efficacy of our method in mitigating biases across various large-scale pre-trained models. This further validates the versatility and applicability of our approach in addressing bias within the realm of natural language processing models, underscoring its potential utility in enhancing the fairness and impartiality of machine learning applications.
A7.4 Parameter Search Experiments
Table A4: The experimental results of parameter exploration. Unlisted values are identical to those of the base model.
Parameters Performance Fairness
$\gamma$ $\alpha$ $\sigma$ Acc $\uparrow$ F1-macro $\uparrow$ Gender $\downarrow$ Age $\downarrow$ Country $\downarrow$ Ethnicity $\downarrow$ AAB $\downarrow$
base 0.1 0.1 0.1 83.94 80.31 2.79 6.21 3.61 4.84 4.36
(a) 0.05 84.20 80.60 2.90 6.15 4.31 5.00 4.59
0.15 85.12 81.74 3.90 7.58 4.43 4.68 5.15
[…]
0.15 85.25 82.00 3.51 7.53 4.66 4.53 5.06
[…]
0.15 83.92 80.32 2.55 6.27 3.98 4.94 4.43
[…]
0.05 0.05 85.09 81.30 3.06 6.83 4.33 4.53 4.69
0.05 0.05 0.15 85.65 82.46 3.77 7.77 5.36 4.70 5.40
[…]
0.05 0.05 85.36 81.81 4.19 7.76 4.53 5.06 5.38
0.15 0.05 0.05 82.85 79.54 3.60 6.99 4.39 3.98 4.74
[…]
0.05 0.05 84.87 81.29 3.54 6.66 4.31 4.83 4.83
0.05 0.15 0.05 84.82 81.50 4.08 7.75 4.92 4.61 5.34
[…]
0.15 0.15 84.53 81.48 4.02 8.04 5.08 4.11 5.31
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
[…]
0.15 0.15 84.94 81.57 3.83 7.24 4.04 4.81 4.98
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
[…]
0.15 0.15 86.15 82.84 3.99 8.33 5.27 4.75 5.58
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
Table A5: The detailed experimental results of debiasing preservation capability. The columns and rows of the table are the methods and abbreviations for each task sequence respectively.
Sequence CLF REPINA IDBR DKT RID
aceg -3.22 -0.01 -1.55 -1.27 -1.66
acge -0.64 0.05 0.26 0.05 -2.70
aecg -0.52 0.03 -1.41 -1.45 -2.46
aegc -1.91 0.00 -1.68 -0.81 -0.67
agce -3.03 0.03 -2.55 -0.84 -1.63
agec -4.18 0.00 -0.15 0.23 0.42
caeg -0.23 -0.08 -0.72 -2.94 -1.58
cage -3.07 0.04 -1.57 -1.67 -2.46
ceag -1.85 -0.03 -0.57 -1.33 -2.00
cega -1.45 0.00 0.09 0.76 -1.82
cgae -3.29 0.06 0.64 -1.43 -0.41
cgea -0.43 0.02 0.17 0.31 0.47
eacg -2.25 -0.06 -1.68 -2.81 -3.03
eagc -1.18 -0.03 -3.14 -2.17 -2.35
ecag 0.05 -0.04 -0.72 -1.40 -1.12
ecga -1.28 0.00 -0.67 -1.11 -0.97
egac -4.64 -0.03 -2.35 -1.41 0.76
egca -3.17 -0.05 -0.51 -0.36 -1.01
gace -1.52 0.05 -0.86 -1.70 -3.40
gaec -0.40 0.02 0.24 0.20 0.31
gcae -0.36 0.07 0.79 -0.79 0.06
gcea 0.99 -0.02 0.71 0.27 -0.40
geac -0.35 0.03 1.77 0.60 0.02
geca -0.52 -0.01 0.72 1.32 1.22
Avg -1.60 0.00 -0.61 -0.82 -1.10
We conducted exploratory experiments on different parameters, which are shown in Table A4. Each parameter has a unique influence on the model’s performance and fairness. Isolating the parameters helps to understand their individual effects, but the best setting would likely depend on the specific performance and fairness goals of the study or the application in which the model is to be used:
Impact of $\gamma$ Change: Experiment (a), (e) and (i) show how varying $\gamma$ influences outcomes. With $\alpha$ and $\sigma$ fixed at 0.1, increasing $\gamma$ (from base to experiment a) reduces performance slightly but improves fairness in gender. With $\alpha$ and $\sigma$ at 0.05, increasing $\gamma$ (experiment e) seems to balance both performance and fairness reasonably well. Experiment (i) also increases $\gamma$, but with $\alpha$ and $\sigma$ at 0.15, showing improvements in performance and fairness, except for age and country.
Impact of $\alpha$ Change: Experiment (b), (f) and (h) all tweak $\alpha$. In experiment (b) with $\gamma$ and $\sigma$ at 0.1, decreasing $\alpha$ slightly decreases performance metrics but improves fairness. With $\gamma$ and $\sigma$ at 0.05, the change in $\alpha$ does not significantly impact performance but seems to worsen fairness in age, country, and ethnicity.
Impact of $\sigma$ Change: Experiment (c), (d) and (g) vary $\sigma$ with $\gamma$ and $\alpha$ at 0.05, 0.1 and 1.05 respectively. In these cases, decreasing $\sigma$ tends to improve fairness in gender and ethnicity but at a slight cost to performance. Experiment (g) increases $\sigma$ with $\gamma$ and $\alpha$ at 0.15, showing an improvement in performance but worsening fairness metrics, notably in gender, country, and ethnicity.
Optimal Settings: Experiment (base) appears to find a middle ground, where increasing $\gamma$ to 0.1 with $\alpha$ and $\sigma$ also at 0.1 offers a balance of improved performance and fairness.
A7.5 Detail of Debiasing Preservation Capability Results
The debiasing preservation capability of different methods is critical to ensure that debiasing performance is maintained across successive tasks without significant degradation. We chose CL as the debiasing method and the results of various methods in preserving debiasing effects over training are presented in Table A5. The results of the debiasing preservation capability evaluation highlight the superior performance of the CLF method in maintaining debiasing effects over time. This method’s unique approach of integrating contrastive learning and continual learning frameworks helps in effectively mitigating biases while preserving overall model performance. Other methods like REPINA show no improvement in bias preservation, while methods like IDBR and TKD exhibit moderate improvements. The significant improvements observed with the CLF method suggest that it is a viable solution for continuous debiasing tasks, promoting fairness and reducing biases in pre-trained language models in the long term.
A7.6 Statistical Significance Test
Table A6: Results of Wilcoxon signed-rank test significance analysis. A p value below 0.05 signifies that our method demonstrates statistically significant improvement over the comparison method.
@@
We have conducted a statistical significance test to validate the significance of the experiment results. We performed a Wilcoxon signed-rank test significance analysis [60] on the experimental results of our model CLF (on 24 different debiasing sequences) as well as the experimental results of other continual debiasing methods. We chose ATC and CL as the debiasing methods for significance analysis. The experimental results are shown in Table A6. The p values of the significance analysis of our proposed model and the existing model are shown in the table. Experimental results show that our improvement is significant.

### v2

@@ V-G Experiments on Other Pre-trained Models
In addition to conducting experiments on the BERT model, we further perform our method with other pre-trained language models Deberta-v3-base [54] and RoBERTa-base [55]. The experimental results are shown in the Table IX. Experimental results verify the efficacy of our method in mitigating biases across various large-scale pre-trained models. This further validates the versatility and applicability of our approach in addressing bias within the realm of natural language processing models, underscoring its potential utility in enhancing the fairness and impartiality of machine learning applications.
V-H Statistical Significance Test
TABLE X: Results of Wilcoxon signed-rank test significance analysis. A p value below 0.05 signifies that our method demonstrates statistically significant improvement over the comparison method.
@@ V-H Statistical Significance Test
We have conducted a statistical significance test to validate the significance of the experiment results. We performed a Wilcoxon signed-rank test significance analysis [56] on the experimental results of our model CLF (on 24 different debiasing sequences) as well as the experimental results of other continual debiasing methods. We chose ATC and CL as the debiasing methods for significance analysis. The experimental results are shown in Table X. The p values of the significance analysis of our proposed model and the existing model are shown in the table. Experimental results show that our improvement is significant.
VI Conclusion
In this paper, we focus on the continuous debiasing task, addressing the problem of debiasing in hate speech scenarios with continuously varying bias types. We propose a dataset, HateDebias, to benchmark continuous debiasing tasks. The HateDebias dataset contains 23,276 samples with four bias attributes, corresponding to four sub-datasets. Based on HateDebias, we also propose a continuous-learning-based framework to eliminate emerging biases while preserving model performance in downstream tasks. We evaluate the debiasing performance of our methods and baselines on HateDebias, which demonstrates that our methods yield a promising performance for improving continuous debiasing tasks. In the future, we will extend the CLF framework to handle intersecting bias attributes, which involves developing more sophisticated representation disentanglement mechanisms to isolate and debias multiple overlapping sensitive attributes simultaneously, as well as designing new evaluation metrics to assess fairness across intersecting groups. Additionally, we aim to explore how to avoid unintended bias amplification for non-targeted groups when debiasing intersecting attributes, further enhancing the framework’s applicability to complex fairness challenges.
Acknowledgements
Our work is supported by the National Social Science Fund of China (No. 22BTQ045).